Circular No. 7235

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990806 F. Frontera, Istituto TESRE, CNR, Bologna and University of Ferrara;

M. Capalbi and M. R. Daniele, BeppoSAX SDC, Rome; E. Montanari, University of Ferrara; A. Paolino and A. Tesseri, BeppoSAX SOC, Rome; C. Pastor, SAX-OCC, Rome; and G. Gandolfi, IAS, CNR, Rome, on behalf of a larger collaboration, report: “The BeppoSAX Wide Field Camera (WFC) error box of GRB 990806 (α = 3^{h}10^{m}.6, δ = −68^{o}07′, error radius = 3′; GCN 392) was observed, beginning about 8 hours after the burst, with the BeppoSAX Narrow Field Instruments in the period Aug. 6.9267–8.6563 UT. Preliminary analysis of MECS2+3 data shows a previously unknown Xray source, 1SAX J0310.6–6806, located at α = 3^{h}10^{m}35^{s}, δ = −68^{o}06′35′′ (equinox 2000.0), with an error radius of 1′. This position is very close to the centroid of the WFC error box. During the observation, the 1.4–10keV source flux decreased from (8.6 ± 2.4) × 10^{3} cts s^{−1}, corresponding to (5.5±1.5)×10^{−13} erg s^{−1} cm^{−2}, to (2.3±0.6)×10^{−3} cts s^{−1}, corresponding to (1.5 ± 0.6) × 10^{−13} erg s^{−1} cm^{−2}. For comparison, a ROSAT source, 1RXS J031250.0–680915, out of the WFC error box but visible in the same MECS image, does not show any statistically significant variation during the entire observation. We conclude that 1SAX J0310.6–6806 is likely the x-ray afterglow of GRB 990806.”

CATALOGUE OF COMETARY ORBITS 1999 The thirteenth edition of the Catalogue of Cometary Orbits, complete

through the 1999 July 28 batch of Minor Planet Circulars, is now available. It contains 1722 orbits for 1688 cometary apparitions referring to 1036 different comets. A tabulation provides osculating elements for 133 of the 140 numbered periodic comets at the standard epochs 1999 Aug. 10, 2000 Feb. 26, 2000 Sept. 13 and 2001 Apr. 1. In a departure from previous practice, the single-apparition comets with calculated periods of less than 200 years are now divided into two groups, with an arbitrary boundary at 30 years separating them into 58 short-period comets and 13 intermediate-period comets, the latter group henceforth retaining the C/ prefixes they would have had when only parabolic orbits were available, rather than the D/ prefixes that have recently been causing undesirable confusion in connection with three recent comets. ‘Original’ and ‘future’ reciprocal semimajor axes are given for 330 high-quality long-period orbits. The 127-page Catalogue sells for $30.00 ($40.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $50.00.

1999 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Gareth V. Williams